OpenStudy (anonymous):

a sample of a radioactive substance was put in storage 6 years ago. It's half-life is 5.5 years. If 8.3 gm of the original substance remains in the sample now, what is the initial quantity of the substance (accurate to two significant figures)?

OpenStudy (anonymous):

so your equation that you have to know and work with is...\[(\frac{1}{2})^{\frac{ 6 }{ 5.5 }} = \frac{8.3}{x}\] with x being your orignial amount

OpenStudy (anonymous):

that fraction on top is \[\frac{6}{5.5}\] btw

OpenStudy (anonymous):

that should help you out a lot